# Partial Renormalization of Quasiparticle Interactions

### v1

@@ I Introduction
The primary challenge in nonlocal QFTs, known as the “vertex problem,” stems from the difficulty in systematically resumming the full nonlocal interaction vertex within the framework of many-body perturbation theory. This resummation process, known as the renormalization of interactions, is crucial for obtaining a well-behaved perturbative expansion that accurately captures the low-energy physics. Existing quantum many-body approaches, such as Dyson-Schwinger equations Dyson (1949); Schwinger (1951), functional renormalization group equations (fRG)Wetterich (1993); Morris (1994); Metzner et al. (2012), and parquet equationsLi et al. (2016); Eckhardt et al. (2018); Valli et al. (2015), attempt to tackle this challenge by employing a skeleton diagrammatic framework to resum the full interaction vertex. However, despite their successes in various applications, these methods encounter significant obstacles that lie at the heart of the vertex problem.
@@
Our scheme goes beyond a mere technical development; it generalizes the well-established field-theoretical renormalization scheme Peskin (2018); Hepp (1966); Zimmermann (1968); Zlmmbrmann (2000), originally developed for local QFTs in high-energy physics Bethe (1947); Schwinger (1948); Feynman (1948), to nonlocal QFTs in many-body physics. The cornerstone of our approach is a momentum-resolved renormalization condition based on projection operators, which extends the concept of coupling constant renormalization to nonlocal interactions. This generalized condition gives rise to renormalized Feynman diagrams and counterterms that are functionals of the nonlocal interactions. To systematically derive these diagrams, we introduce a parquet algorithm that efficiently constructs high-order vertex diagrams by combining lower-order subdiagrams. Moreover, we develop a dedicated diagrammatic Monte Carlo algorithm (DiagMC) Prokof’ev and Svistunov (1998, 2008); Kozik et al. (2010); Rossi (2017); Prokof’ev and Svistunov (2007); Carlström (2021); Taheridehkordi et al. (2019) specifically designed to efficiently compute the resulting diagrammatic series. By generalizing the field-theoretical renormalization scheme to nonlocal QFTs, our approach opens up an alternative avenue for addressing the vertex problem.
@@ II.2 Field-theoretical Renormalization Scheme
To develop a framework for solving nonlocal QFTs, it is instructive to review the field-theoretical renormalization scheme Peskin (2018); Hepp (1966); Zimmermann (1968); Zlmmbrmann (2000), which has been remarkably successful in tackling local QFTs such as quantum electrodynamics (QED) Bethe (1947); Schwinger (1948); Feynman (1948). In local QFTs, bare perturbation theory often proves inadequate due to the presence of large parameters arising from the renormalization of interactions. These large parameters typically manifest as ultraviolet (UV) divergences in loop integrals, rendering the perturbative expansion ill-defined. The field-theoretical renormalization scheme systematically addresses this issue by introducing counterterms in the action. These counterterms are determined by imposing renormalization conditions that establish a well-defined relationship between the bare and renormalized quantities at a specific energy scale, known as the renormalization scale. Although condensed matter systems generally do not suffer from UV divergences due to the presence of natural cutoffs such as the lattice spacing, the challenge of dealing with large parameters arising from the renormalization of interactions remains pertinent.
@@
In this work, we explore a renormalization scheme of nonlocal interactions with the following strategy: 1) Developing an ansatz for the renormalized coupling $R$ that accurately captures its dependence on momentum and frequency. This task requires going beyond the standard local QFT framework and incorporating the specific features of quantum many-body systems, such as the Fermi surface and the associated length and energy scales. 2) Adapting the standard renormalized diagrammatic technique of local QFT to nonlocal theories. This requires development of efficient algorithms to derive and compute the diagrams as well as the counterterms.
@@
Figure 1: a) The projection operator (the red box) to extract the forward-scattering part of the particle-hole-reducible vertex function. The incoming momenta/frequencies $k_{1}$ and $k_{2}$ are projected onto the Fermi surface ($k=({\mathbf{k}},i\omega_{n})\rightarrow\bar{k}=(k_{F}{\mathbf{k}}/|{\mathbf{k}}|,i\omega_{0})$), and two outgoing momenta/frequencies are projected to the EXACT forward-scattering direction. b) Examples of the third-order renormalized diagrams of the vertex function $\Gamma_{4}$. It demonstrates the cancellation between the counter-diagrams (in the red box) and the corresponding vertex subdiagrams.
@@ III Projection-based Renormalization Condition
We focus on the 4-point vertex function, $\Gamma_{4}(k_{1},k_{2};k^{\prime}_{1},k^{\prime}_{2})$, which depends on four momentum-frequency vectors: two incoming particles with $k_{1}$ and $k_{2},i$, and two outgoing particles with $k^{\prime}_{1}$ and $k^{\prime}_{2}$. Due to the momentum-frequency conservation law, $k_{1}+k_{2}=k^{\prime}_{1}+k^{\prime}_{2}$, only three of these four vectors are independent. The 4-point vertex function $\Gamma_{4}$ consists of two parts: the bare interaction $u_{k_{1}k_{2};k^{\prime}_{1}k^{\prime}_{2}}$ and the quantum many-body correction $\Delta\Gamma_{4}(k_{1},k_{2};k^{\prime}_{1},k^{\prime}_{2})$. The bare interaction $u$ possesses a simple analytic form and captures the essential UV physics. In contrast, $\Delta\Gamma_{4}$ exhibits a complex dependence on momentum and frequency encodes the emergent infrared (IR) physics.
@@ IV Renormalized Feynman Diagrammatic Expansion
Generating high-order diagrams in the renormalized expansion is a significant challenge. To address this problem, we build upon the algorithm proposed recently in Ref. Hou et al. (2024), which systematically constructs a Feynman diagrammatic expansion using the parquet equations. These equations, in their conventional form, are a set of self-consistent relations that connect the four-point vertex function to its irreducible components in different channels. The algorithm repurposes the parquet equations by expanding them perturbatively, transforming them into a generator that systematically builds high-order four-point vertex diagrams from lower-order subdiagrams. The resulting high-order Feynman diagrams are not represented as individual diagrams; instead, they are organized into a single comprehensive computational graph of nodes and leaves, where the leaves represent the propagators and interactions, and the nodes represent mathematical operations. This computational graph representation allows for the reuse of many subnodes, greatly simplifying the integrand evaluation process compared to the individual diagram representation.
@@
To overcome this limitation, we extend the algorithm by introducing a set of renormalized parquet equations, as depicted in Fig. 2. Unlike the original parquet equations that involve the bare coupling, the renormalized parquet equations operate on the renormalized interactions. We then perturbatively expand the renormalized parquet equations to generate the renormalized diagrammatic expansion. The hallmark of our iterative scheme is the simultaneous generation of all diagrammatic topologies and counter-diagrams (namely, the subdiagrams in the red boxes in Fig.1 and Fig. 2). Specifically, one starts with the first-order physical vertex function $\Gamma_{4}^{1}=R$ and the bare coupling parameterized with the renormalized interaction $u^{1}=R$, and then substitutes them into the renormalized parquet equations to generate the second-order diagrams of the particle-hole channel $\Gamma_{ph}$, the particle-hole-exchanged channel $\tilde{\Gamma}_{ph}$, and the particle-particle channel $\Gamma_{pp}$. This process is repeated until all diagrams up to order $N$ are generated. Notably, this all-in-one scheme is more straightforward than conventional iterative algorithms such as Bogoliubov’s R operation Bogoliubov and Parasiuk (1957); Hepp (1966); Zimmermann (1968); Zlmmbrmann (2000), where one first needs to list all topologies of bare diagrams and then iteratively generates the corresponding counter-diagrams for each topology.
@@ V Imaginary-time Representation
DiagMC algorithms with renormalized expansions have found numerous applications in quantum many-body physics, yielding state-of-the-art results for various systems. For instance, BDMC has been successfully applied to the Fermi polaron problem Prokof’ev and Svistunov (2007); Mishchenko et al. (2014), the Hubbard model Kozik et al. (2010); Deng et al. (2015); Carlström (2023), the resonant Fermi gas Van Houcke et al. (2012); Rossi et al. (2018), and the frustrated spin systems Kulagin et al. (2013a, b); Huang et al. (2016); Wang et al. (2020), providing accurate predictions for the equation of state and correlation functions. The CDet method has been employed to calculate the self-energy and spectral functions in the Hubbard model Rossi (2017); Lenihan et al. (2022); IV et al. (2022); Simkovic et al. (2022) and a variety of other quantum many-body problems Van Houcke et al. (2020), while the combinatorial approach has been used to study the equation of state of the 2D SU(N) Hubbard model in an experimentally relevant regime Kozik (2023). Additionally, the Taylor-mode automatic differentiation approach has been utilized to investigate the electron gas problem with long-range Coulomb interactions, yielding precise values for the quasiparticle effective mass and other low-energy quantities Chen and Haule (2019); Hou et al. (2022); LeBlanc et al. (2022); Hou et al. (2024).
@@
Figure 4: Application of the projection operator (red box) to a particle-hole vertex diagram. The operator projects the two incoming momenta onto the Fermi surface, mapping $k\rightarrow\bar{k}$, as explained in the main text. It also projects the temporal variables of all four external legs to the same value as the left incoming leg, ensuring that all external frequencies are projected to $i\omega_{0}$ in the Matsubara frequency representation. This projection operation is essential for isolating the dominant low-energy contributions to the vertex function. It is important to note that the external momenta experience an abrupt change before and after the projection (across the red box). Notably, the projected vertex function still obeys the momentum conservation law. The application of the projection operator is a crucial step in the renormalization procedure, as it allows for the identification and selective resummation of the most important contributions to the vertex function, while avoiding the computational complexity associated with the full vertex function.
@@ VI Application
where $\theta$ is the angle between the two incoming momenta, and $N_{F}$ is the density of states on the Fermi surface. In the paramagnetic limit, the scattering amplitude is decomposed into a spin-symmetric component $A^{s}$ and a spin-asymmetric component $A^{a}$. Each component can be further decomposed into different angular momentum channels using the Legendre transform $A^{s,a}(\theta)=\sum_{l=0}^{\infty}(2l+1)A^{s,a}_{l}P_{l}(\cos\theta)$, where $A^{s,a}_{l}$ are the Legendre coefficients and $P_{l}(\cos\theta)$ are the Legendre polynomials. The Landau quasiparticle interaction $F^{s,a}(\theta)$, on the other hand, is defined as the 4-vertex function in the zero transfer momentum and small transfer frequency limit, averaged over the Fermi surface. It is related to the scattering amplitude the relation: $A_{l}^{s,a}=F_{l}^{s,a}/\left(1+F_{l}^{s,a}\right)$, where $F_{l}^{s,a}$ are the Legendre coefficients of $F^{s,a}(\theta)$.
@@ VII Conclusion
Acknowledgements – We are grateful to Gabriel Kotliar, Kristjan Haule, Nikolay Prokof’ev, Boris Svistunov, Youjin Deng, Christopher Amey, Yashar Komijani, and John Sous for stimulating discussions, valuable input, and suggestions that greatly improved the readability of this manuscript. We also thank Alessandro Toschi, Piers Coleman and Romain Vasseur for enlightening conversations that helped shape our understanding of the subject. This work was supported by the National Natural Science Foundation of China under Grant No. 12047503 and the Simons Foundation through the Many Electron Collaboration. K.C. acknowledges additional support from the National Natural Science Foundation of China under Grant No. 12047503.

### v2

@@ I Introduction
The primary challenge in nonlocal QFTs, known as the “vertex problem”, stems from the difficulty in systematically resumming the full nonlocal interaction vertex within the framework of many-body perturbation theory. This resummation process, known as the renormalization of interactions, is crucial for obtaining a well-behaved perturbative expansion that accurately captures the low-energy physics. Existing quantum many-body approaches, such as Dyson-Schwinger equations Dyson (1949); Schwinger (1951), functional renormalization group equations (fRG)Wetterich (1993); Morris (1994); Metzner et al. (2012), and parquet equationsLi et al. (2016); Eckhardt et al. (2018); Valli et al. (2015), attempt to tackle this challenge by employing a skeleton diagrammatic framework to resum the full interaction vertex. However, despite their successes in various applications, these methods encounter significant obstacles that lie at the heart of the vertex problem.
@@
Our scheme goes beyond a mere technical development; it generalizes the well-established field-theoretical renormalization scheme Peskin (2018); Hepp (1966); Zimmermann (1968, 2000), originally developed for local QFTs in high-energy physics Bethe (1947); Schwinger (1948); Feynman (1948), to nonlocal QFTs in many-body physics. The cornerstone of our approach is a momentum-resolved renormalization condition based on projection operators, which extends the concept of coupling constant renormalization to nonlocal interactions. This generalized condition gives rise to renormalized Feynman diagrams and counterterms that are functionals of the nonlocal interactions. To systematically derive these diagrams, we introduce a parquet algorithm that efficiently constructs high-order vertex diagrams by combining lower-order subdiagrams. Moreover, we develop a dedicated diagrammatic Monte Carlo algorithm (DiagMC) Prokof’ev and Svistunov (1998, 2008); Kozik et al. (2010); Rossi (2017); Prokof’ev and Svistunov (2007); Carlström (2021); Taheridehkordi et al. (2019) specifically designed to efficiently compute the resulting diagrammatic series. By generalizing the field-theoretical renormalization scheme to nonlocal QFTs, our approach opens up an alternative avenue for addressing the vertex problem.
@@ II.2 Field-theoretical Renormalization Scheme
To develop a framework for solving nonlocal QFTs, it is instructive to review the field-theoretical renormalization scheme Peskin (2018); Hepp (1966); Zimmermann (1968, 2000), which has been remarkably successful in tackling local QFTs such as quantum electrodynamics (QED) Bethe (1947); Schwinger (1948); Feynman (1948). In local QFTs, bare perturbation theory often proves inadequate due to the presence of large parameters arising from the renormalization of interactions. These large parameters typically manifest as ultraviolet (UV) divergences in loop integrals, rendering the perturbative expansion ill-defined. The field-theoretical renormalization scheme systematically addresses this issue by introducing counterterms in the action. These counterterms are determined by imposing renormalization conditions that establish a well-defined relationship between the bare and renormalized quantities at a specific energy scale, known as the renormalization scale. Although condensed matter systems generally do not suffer from UV divergences due to the presence of natural cutoffs such as the lattice spacing, the challenge of dealing with large parameters arising from the renormalization of interactions remains pertinent.
@@
Figure 1: a) The projection operator (the red box) to extract the forward-scattering part of the particle-hole-reducible vertex function. The incoming momenta/frequencies $k_{1}$ and $k_{2}$ are projected onto the Fermi surface ($k=({\mathbf{k}},i\omega_{n})\rightarrow\bar{k}=(k_{F}{\mathbf{k}}/|{\mathbf{k}}|,i\pi T)$), and two outgoing momenta/frequencies are projected to the EXACT forward-scattering direction. b) Examples of the third-order renormalized diagrams of the vertex function $\Gamma_{4}$. It demonstrates the cancellation between the counter-diagrams (in the red box) and the corresponding vertex subdiagrams.
@@ III Projection-based Renormalization Condition
We focus on the 4-point vertex function, $\Gamma_{4}(k_{1},k_{2};k^{\prime}_{1},k^{\prime}_{2})$, which depends on four momentum-frequency vectors: two incoming particles with $k_{1}$ and $k_{2}$, and two outgoing particles with $k^{\prime}_{1}$ and $k^{\prime}_{2}$. Due to the momentum-frequency conservation law, $k_{1}+k_{2}=k^{\prime}_{1}+k^{\prime}_{2}$, only three of these four vectors are independent. The 4-point vertex function $\Gamma_{4}$ consists of two parts: the bare interaction $u_{k_{1}k_{2};k^{\prime}_{1}k^{\prime}_{2}}$ and the quantum many-body correction $\Delta\Gamma_{4}(k_{1},k_{2};k^{\prime}_{1},k^{\prime}_{2})$. The bare interaction $u$ possesses a simple analytic form and captures the essential UV physics. In contrast, $\Delta\Gamma_{4}$ exhibits a complex dependence on momentum and frequency encodes the emergent infrared (IR) physics.
@@ IV Renormalized Feynman Diagrammatic Expansion
Generating high-order diagrams in the renormalized expansion is a significant challenge. To address this problem, we build upon the algorithm proposed recently in Ref. Hou et al. (2024), which systematically constructs a Feynman diagrammatic expansion using the parquet equations. These equations, in their conventional form, are a set of self-consistent relations that connect the four-point vertex function to its irreducible components in different channels. The algorithm repurposes the parquet equations by expanding them perturbatively, transforming them into a generator that systematically builds high-order four-point vertex diagrams from lower-order subdiagrams. The resulting high-order Feynman diagrams are not represented as individual diagrams; instead, they are organized into a single comprehensive computational graph of nodes and leaves, where the leaves represent the propagators and interactions, and the nodes represent mathematical operations. This computational graph representation allows for the reuse of many subnodes, greatly simplifying the integrand evaluation process compared to the individual diagram representation. Ref. Kugler (2018) has utilized a similar idea to count the number of Feynman diagrams of vertex function.
@@
To overcome this limitation, we extend the algorithm by introducing a set of renormalized parquet equations, as depicted in Fig. 2. The equations are derived by re-expanding the bare couplings in the irreducible vertex functions as a power series of renormalized couplings, removing the UV physics in the original bare parquet equations. We then perturbatively expand the renormalized parquet equations to generate the renormalized diagrammatic expansion. The hallmark of our iterative scheme is the simultaneous generation of all diagrammatic topologies and counter-diagrams (namely, the subdiagrams in the red boxes in Fig.1 and Fig. 2). Specifically, one starts with the first-order physical vertex function $\Gamma_{4}^{1}=R$ and the bare coupling parameterized with the renormalized interaction $u^{1}=R$, and then substitutes them into the renormalized parquet equations to generate the second-order diagrams of the particle-hole channel $\Gamma_{ph}$, the particle-hole-exchanged channel $\tilde{\Gamma}_{ph}$, and the particle-particle channel $\Gamma_{pp}$. This process is repeated until all diagrams up to order $N$ are generated. Notably, this all-in-one scheme is more straightforward than conventional iterative algorithms such as Bogoliubov’s R operation Bogoliubov and Parasiuk (1957); Hepp (1966); Zimmermann (1968, 2000), where one first needs to list all topologies of bare diagrams and then iteratively generates the corresponding counter-diagrams for each topology.
@@
The current algorithm implementation generates fully irreducible diagrams in $I$ by exhaustively considering all possible topology at a given order Chen and Haule (2019). For future work, it would be valuable to explore an alternative approach that involves taking the functional derivative of the parquet self-energy diagrams Kugler and von Delft (2018).
@@ V Imaginary-time Representation
DiagMC algorithms with renormalized expansions have found numerous applications in quantum many-body physics, yielding state-of-the-art results for various systems. For instance, BDMC has been successfully applied to the Fermi polaron problem Prokof’ev and Svistunov (2007); Mishchenko et al. (2014), the Hubbard model Kozik et al. (2010); Deng et al. (2015); Carlström (2023), the resonant Fermi gas Van Houcke et al. (2012); Rossi et al. (2018), and the frustrated spin systems Kulagin et al. (2013a, b); Huang et al. (2016); Wang et al. (2020), providing accurate predictions for the equation of state and correlation functions. The CDet method has been employed to calculate the self-energy and spectral functions in the Hubbard model Rossi (2017); Lenihan et al. (2022); IV et al. (2022); Simkovic et al. (2022) and a variety of other quantum many-body problems Van Houcke et al. (2020), while the combinatorial approach has been used to study the equation of state of the 2D SU(N) Hubbard model in an experimentally relevant regime Kozik (2023). Additionally, the Taylor-mode automatic differentiation approach has been utilized to investigate the electron gas problem with long-range Coulomb interactions, yielding precise values for the quasiparticle effective mass and other low-energy quantities Chen and Haule (2019); Haule and Chen (2022); Hou et al. (2022); LeBlanc et al. (2022); Hou et al. (2024).
@@
Figure 4: Application of the projection operator (red box) to a particle-hole vertex diagram. The operator projects the two incoming momenta onto the Fermi surface, mapping $k\rightarrow\bar{k}$, as explained in the main text. It also projects the temporal variables of all four external legs to the same value as the left incoming leg, ensuring that all external frequencies are projected to $\omega_{0}=\pi T$ in the Matsubara frequency representation. This projection operation is essential for isolating the dominant low-energy contributions to the vertex function. It is important to note that the external momenta experience an abrupt change before and after the projection (across the red box). Notably, the projected vertex function still obeys the momentum conservation law. The application of the projection operator is a crucial step in the renormalization procedure, as it allows for the identification and selective resummation of the most important contributions to the vertex function, while avoiding the computational complexity associated with the full vertex function.
@@ VI Application
where $\theta$ is the angle between the two incoming momenta, and $N_{F}$ is the density of states on the Fermi surface. In the paramagnetic limit, the scattering amplitude is decomposed into a spin-symmetric component $A^{s}$ and a spin-asymmetric component $A^{a}$. Each component can be further decomposed into different angular momentum channels using the Legendre transform $A^{s,a}(\theta)=\sum_{l=0}^{\infty}(2l+1)A^{s,a}_{l}P_{l}(\cos\theta)$, where $A^{s,a}_{l}$ are the Legendre coefficients and $P_{l}(\cos\theta)$ are the Legendre polynomials. The Landau quasiparticle interaction $F^{s,a}(\theta)$, on the other hand, is defined as the 4-vertex function in the zero transfer momentum and small transfer frequency limit, averaged over the Fermi surface. It is related to the scattering amplitude by the relation: $A_{l}^{s,a}=F_{l}^{s,a}/\left(1+F_{l}^{s,a}\right)$, where $F_{l}^{s,a}$ are the Legendre coefficients of $F^{s,a}(\theta)$.
@@ VII Conclusion
Acknowledgements – We are grateful to Gabriel Kotliar, Kristjan Haule, Jan von Delft, Fabian Kugler, Nikolay Prokof’ev, Boris Svistunov, Youjin Deng, Christopher Amey, Yashar Komijani, and John Sous for stimulating discussions, valuable input, and suggestions that greatly improved the readability of this manuscript. We also thank Alessandro Toschi, Piers Coleman and Romain Vasseur for enlightening conversations that helped shape our understanding of the subject. This work was supported by the National Natural Science Foundation of China under Grant No. 12047503 and the Simons Foundation through the Many Electron Collaboration. K.C. acknowledges additional support from the National Natural Science Foundation of China under Grant No. 12047503.